OpenStudy (anonymous):

1/4a+8=1/2a solve for a

OpenStudy (texaschic101):

subtract 8 from each side 1/4a = 1/2a - 8 now subtract 1/2a from each side 1/4a - 1/2a = -8 now multiply the entire equation by the LCD which is 4 a - 2a = - 32 combine like terms -a = -32 a = 32 check... 1/4a + 8 = 1/2a 1/4(32) + 8 = 1/2(32) 32/4 + 8 = 32/2 8 + 8 = 16 16 = 16 (correct) a = 32